# DFDT: Dynamic Fast Decision Tree for IoT Data Stream Mining on Edge Devices

### v1

@@ Abstract
The Internet of Things generates massive data streams, with edge computing emerging as a key enabler for online IoT applications and 5G networks. Edge solutions facilitate real-time machine learning inference, but also require continuous adaptation to concept drifts. Ensemble-based solutions improve predictive performance, but incur higher resource consumption, latency, and memory demands. This paper presents DFDT: Dynamic Fast Decision Tree, a novel algorithm designed for energy-efficient memory-constrained data stream mining. DFDT improves hoeffding tree growth efficiency by dynamically adjusting grace periods, tie thresholds, and split evaluations based on incoming data. It incorporates stricter evaluation rules (based on entropy, information gain, and leaf instance count), adaptive expansion modes, and a leaf deactivation mechanism to manage memory, allowing more computation on frequently visited nodes while conserving energy on others. Experiments show that the proposed framework can achieve increased predictive performance (0.43 vs 0.29 ranking) with constrained memory and a fraction of the runtime of VFDT or SVFDT.
Keywords Hoeffding tree  $\cdot$ data stream  $\cdot$ edge computing  $\cdot$ IoT
1 Introduction
@@
The Internet of Things (IoT) connects a vast network of physical devices, generating massive, high-speed data streams. To extract valuable insights from this ever-growing stream of raw sensor data, the edge computing paradigm has emerged as a key enabler for low-latency IoT applications and 5G networks, bridging the gap between cloud services and end-users. However, benefiting from edge solutions not only implies bringing machine learning models for real-time inference, but also continuous updates to adapt to the evolving nature of unbounded streams. Unlike batch learning, where all training data is available upfront, data streams arrive incrementally. Thus, making models susceptible to becoming outdated, especially due to concept drifts, i.e., changes in data distribution over time [6].
@@
To circumvent these challenges, many stream mining algorithms have been proposed, with tree-based methods being the state-of-the-art for tabular data sources. Their success can be attributed to approximation-based splitting [4], with incrementally updated statistical summaries of entropy-based metrics, e.g. information gain, to determine whether the observed utility of a split is statistically significant when the data distribution is unknown, e.g. via the Hoeffding bound (HB). To find the split value, for categorical attributes, one can simply do a split for each category, and group them by similarity criteria [4], while for continuous attributes, one can compute the Gaussian for each combination with the target classes and select the highest posterior probability [5].
Moreover, combining these into ensemble strategies is another effective approach. Upon a concept drift, the diversity among trees can be exploited as some local minima are less affected than others. Affected components can be dynamically deleted, added, and combined with a voting scheme, weighting different weak learners based on their past performance [10, 11]. However, with data stream researchers shifting their focus to ensemble-based solutions prioritizing increased predictive performance, these come at the cost of higher resource consumption, increased latency, and greater memory requirements. This trade-off is particularly problematic in memory-scarce edge environments, making it essential to revisit strategies that allow to control the tree growth efficiency, by delaying the expansion of less confident branches.
To address this challenge, a novel algorithm tailored for memory-constrained data stream mining is introduced, coined as DFDT: Dynamic Fast Decision Tree, which consolidates several key contributions from prior research. The algorithm employs dynamic, adaptive strategies for controlling decision tree growth, specifically by adjusting the grace period and tie thresholds during the tree building process, according to the incoming data samples. Moreover, the algorithm incorporates a stricter set of rules based on entropy, information gain, and the number of instances observed at the leaf. To further manage memory constraints, a deactivation mechanism is introduced for low activity leaves, so that the algorithm can grow more organically depending on the distribution and number of instances observed at each leaf of the tree. Thus allowing to spend more computation on the most visited nodes while saving energy on others.
Experiments conducted on various benchmark datasets, evaluating the accuracy, memory usage, and runtime of DFDT, along with statistical tests to assess significant differences, demonstrate that the proposed adaptive strategies achieve enhanced efficiency (ranking 0.74 vs. 0.65). An ablation study of the algorithm’s key components further highlights that omitting adaptive expansion modes with leaf deactivation seems to lead to even superior overall efficiency, by increasing accuracy and reducing computation time, at the cost of memory. The paper is structured as follows: Section 2 reviews contributions from prior research that underpin the proposed algorithm. Section 3 introduces DFDT, detailing its core concepts and pseudocode. Section 4 presents a comprehensive ablation study of the algorithm components, and compares the performance of DFDT with VFDT, svfdt_I, and svfdt_II. Finally, Section 5 outlines conclusions and future research directions.
2 Related work
@@
The core component for incrementally constructing decision trees (DTs) on edge devices, while maintaining a fixed time complexity per sample, is approximation-based splitting [4]. As new instances arrive, they are processed from the root to a leaf node, updating statistics at each node. These updates are used to periodically adjust heuristic values for each attribute, such as information gain (IG) and Gini index (GI) [4]. The tree attempts to perform splits based on a statistical bound that quantifies the confidence interval for the heuristic function, given a minimum amount of data. Typically, DTs compare the Hoeffding bound (HB) to the difference in evaluation between the best and second-best attribute splits [4]. When this difference exceeds the bound, the leaf node is split into child nodes.
@@
Rules. While traditional incremental DTs exclusively focus on evaluating the top two attributes, one can introduce extra splitting rules to promote even more valuable splits. For example, one can incorporate the fluctuation of the HB, tracking its mean, minimum, and maximum values, along with an accuracy metric, as a pre-pruning condition for split decisions [17]. Alternatively, constraints can be applied that compare current metrics against historical data and cross-leaf information, while introducing a skipping mechanism to bypass these constraints when significant changes in the data are detected [3]. Furthermore, a constraint can be added to assess whether the best-ranked feature, chosen for splitting at a leaf node, provides substantial gains compared to the gains observed in previous splits within the same branch [1].
@@
Tie break. A critical aspect of incremental DTs is the tie-breaking procedure. While using the statistical difference in IG between two attributes helps control tree growth, competition between two equally favored split candidates can hinder progress, especially when either option would be equally suitable. To mitigate this, if the difference in heuristic values exceeds a predefined tie threshold, denoted as $\tau$, the split is performed. This threshold effectively controls the minimum rate of tree growth, with the attributes’ ability to separate before reaching the threshold influencing the speed at which the tree expands. However, a fixed $\tau$ value may cause ties to be broken prematurely, before a meaningful decision can be made, due to a lack of suitable candidates rather than a true tie situation. To address this, alternative approaches can be employed, such as comparing the difference between the best and second-best candidates with the difference between the second-best and worst candidates [12], or designing an adaptive tie threshold that is dynamically calculated from the mean of HB, which has been shown to be proportionally related to the input stream samples [16].
@@
Grace period. To prevent premature splits with small sample sizes that undermine the validity of the HB, a grace period $n_{\min}$ specifies the minimum number of instances a leaf must observe before considering a split. However, without leveraging any information from the processed data, this can still result in computationally expensive split attempts or unnecessary delays in predictions. An alternative approach involves detecting frequent tie-breaking situations and applying steps to reduce the frequency of ties. For example, the default tie-breaking wait period can be adjusted by increasing the wait period for the child nodes after each tie is broken. This effect is cumulative until a valid split is found, after which the wait period is reset to the default value [12]. Additionally, local statistics can be used to predict the optimal split time, minimizing delays and unnecessary split attempts. For instance, class distributions from previous split attempts can be used to estimate the minimum number of examples required before the HB is met [13]. Alternatively, $n_{\min}$ can be adapted after an unsuccessful split attempt, adjusting based on the reasons for failure to ensure a split in the next iteration [8]. For example, if the best attributes are not too similar, but their GI difference is insufficient to trigger a split, the solution is to wait for additional examples until the HB decreases enough to be smaller than the GI difference, adjusting the $n_{\min}$ accordingly. Similarly, if the top attributes are very similar in terms of GI, but the tie threshold $\tau$ is not exceeded, more instances are needed to allow the HB to decrease below $\tau$, with $n_{\min}$ adjusted based on their IG difference [8].
@@
Leaf activity. Another memory-efficient strategy is to condition split decisions based on leaf activity. One approach involves comparing the number of samples that fall into existing leaves against those that do not [17]. As more samples fall into existing leaves, one can refrain from updating the tree structure. Alternatively, a normalized measure can be devised to quantify how many instances have been observed at a particular leaf relative to the average number of instances per leaf since its creation [9]. Based on this value, nodes can either be deactivated when their activity is low, halting further splits to conserve resources, or more aggressive expansion strategies can be applied when activity exceeds a threshold, accelerating the growth of important nodes by relaxing the splitting rule to a less strict alternative [9]. These adaptive expansion modes offer a more nuanced approach to tree growth, ensuring computational resources are focused on expanding nodes that significantly influence the model’s accuracy, while avoiding unnecessary splits in less relevant parts of the tree. Alternatively, the least promising leaves can be deactivated based on the probability that examples will reach those leaves and their observed error rate [4]. If a better split is found, the leaves of that split are deactivated, and their statistics are preserved. A new split is then performed, creating new leaves. If, during split re-evaluation, a previously deactivated split is found to be the best option, the saved statistics are restored rather than starting the process from scratch.
DFDT. The proposed algorithm combines the aforementioned four mechanisms to optimize the trade-off between accuracy and resource efficiency. It incorporates dynamically adjustable grace periods [8] and tie thresholds [16, 17], allowing the algorithm to either delay or accelerate splits depending on data variability. Additionally, DFDT combines the adaptive expansion modes [9], determined by leaf activity with stricter and looser splitting conditions [3, 1]. For nodes with low activity, DFDT deactivates the leaf node, conserving memory and computational resources. For nodes with moderate activity, DFDT applies the conservative splitting constraints, ensuring controlled growth. For highly active nodes, DFDT potentially employs a skipping mechanism, enabling rapid growth in response to significant data changes.
3 Methodology
@@
The proposed DFDT integrates previous research advancements to optimize the trade-off between accuracy and resource efficiency of the original VFDT. These modifications are here described in detail, while referring to the pseudo-code in Algorithm 1. The initialization phase creates the root node structure of the decision tree (Alg. 1, Step 2), sets up estimators (Alg. 1, Step 3), and instance counters (Alg. 1, Step 4). This initialization involves only constant-time operations, resulting in a computational complexity of $O(1)$, independent of the number of instances or features. The main loop iterates over each instance $(X,y)$ in the data stream $S$, running $N$ times in total. For each instance, several steps are performed. First, the instance is routed through the tree to the corresponding leaf node, where the prediction $\hat{y}$ is obtained (Alg. 1, Step 6). Assuming a balanced tree, the depth of traversal is $O(\log_{B}|LH|)$, where $|LH|$ is the number of leaf nodes and $B$ is the branching factor. The prediction at the leaf node is a constant-time operation, $O(1)$. Following this, the instance count and feature estimators are updated at the leaf node (Alg. 1, Steps 7, 8). This update takes $O(F)$ time.
@@
1:procedure DFDT($S$: stream, $\delta$: confidence level)
2:     $DFDT\leftarrow$ root and $LH\leftarrow\text{hash of leaves}$
3:     Initialize $H_{LH_{\text{stat}}},n_{\text{stat}},H_{\text{stat}},G_{\text{stat}},HB_{\text{stat}},n_{\text{min}}$
4:     $n,n_{\text{l}},n_{\text{check}_{\text{l}}},n_{\text{leaf}_{\text{l}}},n_{\text{tree}_{\text{l}}}\leftarrow 0$ where $l=\text{root}$
5:     for each $(X,y)$ in $S$ do
6:         Route $(X,y)$ to leaf $l$ and obtain prediction $\hat{y}$
7:         Update feature estimators, and class distribution at $l$ with $(X,y)$
8:         $n\leftarrow n+1$ and $n_{l}\leftarrow n_{l}+1$
9:         if $\frac{(n_{l}-n_{\text{leaf}_{l}})\times|LH|}{n-n_{\text{tree}_{l}}}<0.2$ then
10:              Deactivate $l$
11:         else if $>2$ then
12:              GrowFast $\leftarrow$ True
13:         end if
14:         if (class distribution at $l$ is impure) and ($n_{l}-n_{\text{check}_{l}}>n_{\text{min}}$then
15:              Update $HB_{stat}$ with updated $\epsilon$
16:              if CanSplit(GrowFast, $G(\cdot)$, $\epsilon$, $HB_{stat}$, $H_{stat}$, $G_{stat}$, $n_{stat}$then
17:                  Replace leaf $l$ with a split node, and remove from set of leaves $LH$
18:                  for each leaf branch $b$ of the split do
19:                       Initialize estimators, set post-split distribution at $b$, add to $LH$
20:                       $n_{b},n_{\text{leaf}_{b}},n_{\text{check}_{b}}\leftarrow$ sum of class distribution at $b$ and $n_{\text{tree}_{b}}\leftarrow n$
21:                  end for
22:              else
23:                  $n_{\text{check}_{l}}\leftarrow n_{l}$
24:                  if $\Delta G<\epsilon$ and $\Delta G>\text{avg}(HB_{\text{stat}})$ then
@@
25:                       $n_{\text{min}}\leftarrow\left\lceil\frac{R^{2}\ln(1/\delta)}{2(\Delta G)^{2}}\right\rceil$
26:                  else if $\Delta G<\text{avg}(HB_{\text{stat}})$ and $\epsilon>\text{avg}(HB_{\text{stat}})$ then
@@
27:                       $n_{\text{min}}\leftarrow\left\lceil\frac{R^{2}\ln(1/\delta)}{2(\text{avg}(HB_{\text{stat}}))^{2}}\right\rceil$
28:                  end if
29:              end if
30:         end if
31:     end for
32:end procedure
@@
While the HB ensures that splits are made with statistical confidence, it treats all nodes equally in terms of expansion, despite the fact that not all nodes contribute equally to the decision-making process. To improve VFDT’s efficiency in allocating computational resources, a notion of relative importance can be introduced to dynamically adjust the rate of expansion at each node. To assess the activity level at each node, a fraction parameter is calculated as [9]:
$\text{fraction}=\frac{(n_{l}-n_{\text{leaf}_{l}})\times|LH|}{n-n_{\text{tree}_{l}}}$ (1)
@@
where $(n_{l}-n_{\text{leaf}_{l}})$ is the number of instances observed at a particular leaf $l$ since its creation, $n-n_{\text{tree}_{l}}$ the total instances observed by the tree since the creation of leaf $l$, and $|LH|$ the current total number of leaves. If the fraction is below 0.2, the leaf node is deactivated, halting further splits to save on resources (Alg. 1, Steps 10), which is a constant-time operation, $O(1)$. If the metric falls above 2, a boolean saves the intention to apply a more aggressive expansion strategy, accelerating the growth of important nodes (Alg. 1, Step 12).
@@
To avoid costly evaluations, the algorithm only checks for potential splits once a leaf node accumulates $n_{\text{min}}$ instances since the last splitting opportunity (Alg. 1, Step 14). Then a split attempt is performed (Alg. 1, Step 16), with its pseudocode detailed in Algorithm 2. This check involves computing $G(\cdot)$ values, requiring $O(F)$ time, and sorting the $G(\cdot)$ values (Alg. 2, Step 2), yielding a time complexity of $O(F\log F)$. Next, the algorithm verifies whether the HB or tie-breaking threshold is satisfied (Alg. 2, Step 4). While in the VFDT, the tie threshold ($\tau$), which effectively controls minimum growth speed, remains static, DFDT, instead, dynamically calculates the tie threshold as [16]:
$\tau=\frac{\sum_{i=1}^{k}\epsilon_{i}}{k}$ (2)
where $\epsilon_{i}$ is the HB value calculated at each instance $i$, for $k$ instances. This reduces the need for trial-and-error in selecting a fixed threshold, optimizing tree performance in dynamic environments. As the data stream becomes noisier, the adaptive $\tau$ stabilizes the model. If the HB is satisfied, DFDT utilizes adaptive expansion modes, determined by the fraction-determined boolean. For nodes with moderate activity, DFDT only applies four conservative splitting constraints [3], ensuring controlled growth (Alg. 2, Step 12-15). For highly active nodes, i.e., when the fraction value exceeds 2, DFDT also checks two skipping conditions (Alg. 2, Step 5-11).
@@
1:procedure CanSplit(GrowFast = False, $G(\cdot),\epsilon,HB_{stat},H_{stat},G_{stat},n_{stat}$),
[…]
3:     $G_{\text{best}}=\max(rank)$, $G_{\text{second best}}=\max(rank\setminus\{G_{\text{best}}\})$
4:     if $G_{\text{best}}-G_{\text{second best}}\geq\epsilon$ or $\epsilon<\text{avg}(HB_{\text{stat}})$ then
@@
6:              $C1\leftarrow H_{l}\geq\text{avg}(H_{\text{stat}})+\sigma(H_{\text{stat}})$
7:              $C2\leftarrow G_{\text{best}}\geq\text{avg}(G_{\text{stat}})+\sigma(G_{\text{stat}})$
8:              if C1 $and$ C2 then

### v2

@@ Abstract
The Internet of Things generates massive data streams, with edge computing emerging as a key enabler for online IoT applications and 5G networks. Edge solutions facilitate real-time machine learning inference, but also require continuous adaptation to concept drifts. While extensions of the Very Fast Decision Tree (VFDT) remain state-of-the-art for tabular stream mining, their unregulated growth limit efficiency, particularly in ensemble settings where post-pruning at the individual tree level is seldom applied. This paper presents DFDT, a novel memory-constrained algorithm for online learning. DFDT employs activity-aware pre-pruning, dynamically adjusting splitting criteria based on leaf node activity: low-activity nodes are deactivated to conserve resources, moderately active nodes split under stricter conditions, and highly active nodes leverage a skipping mechanism for accelerated growth. Additionally, adaptive grace periods and tie thresholds allow DFDT to modulate splitting decisions based on observed data variability, enhancing the accuracy–memory–runtime trade-off while minimizing the need for hyperparameter tuning. An ablation study reveals three DFDT variants suited to different resource profiles. Fully compatible with existing ensemble frameworks, DFDT provides a drop-in alternative to standard VFDT-based learners.
Introduction
@@
The Internet of Things (IoT) connects a vast network of physical devices, generating massive, high-speed data streams (gaber2014data). To extract valuable insights from this ever-growing stream, the edge computing paradigm has emerged as a key enabler for low-latency IoT applications and 5G networks, bridging the gap between cloud services and end-users (lourencco2025device). However, benefiting from edge solutions not only implies bringing machine learning models for real-time inference, but also continuous updates to adapt to the evolving nature of unbounded streams. Unlike batch learning, where all training data is available upfront, data streams arrive incrementally (gama2010knowledge; lourencco2025context). Thus, making models susceptible to becoming outdated, especially due to concept drifts, i.e., changes in data distribution over time.
Figure 1: Accuracy x Memory
@@
To circumvent these challenges, many stream mining algorithms have been proposed, with extensions of the Very Fast Decision Tree (VFDT) (domingos2000mining) being the state-of-the-art for tabular data sources. Their success can be attributed to approximation-based splitting, with incrementally updated statistical summaries of entropy-based metrics, e.g. information gain, to determine whether the observed utility of a split is statistically significant when the data distribution is unknown, e.g. via the Hoeffding bound (HB). This enables trees to grow incrementally, refining their partitions of the input space as more data arrives. However, uncontrolled tree growth can lead to excessive memory usage and reduced adaptability.
To counter this, adaptive tree algorithms introduce post-pruning mechanisms for forgetting. For instance, CVFDT (hulten2001mining) evaluates alternate subtrees in parallel, replacing them when they outperform their predecessors. HAT (bifet2009adaptive) generalizes this with multiple recursive replacements, while EFDT (manapragada2018extremely) aggressively prunes subtrees when better local structures emerge. Other methods, like UFFT (gama2005learning) and OnlineTree2 (nunez2007learning), use a short-term memory to stage and evaluate candidate subtrees, enabling smoother transitions and compact models.
Despite these innovations, single decision trees often suffer from instability and limited predictive performance. As a result, the community has largely shifted towards ensemble-based methods, where multiple diverse trees are combined to enhance robustness and adaptability (gomes2017survey; krawczyk2017ensemble). Ensembles handle concept drift at a higher level, dynamically updating individual components based on performance signals and using drift detectors to add, remove, or reweight learners (neves2025online). Thus, reducing the need for post-pruning at the individual tree level. As a result, most state-of-the-art ensembles still adopt the original VFDT as their base learner, often leading to inefficient memory use and redundant computation. Since post-pruning is rarely applied in ensemble settings, this raises a critical question: can pre-pruning strategies be leveraged to proactively regulate VFDT’s growth, preventing unnecessary expansion in uncertain or low-utility regions of the input space?
To this end, a novel algorithm tailored for memory-constrained data stream mining is introduced, coined as DFDT: Dynamic Fast Decision Tree, whose main contributions are:
  • different splitting conditions according to leaf activity, so that the algorithm can grow more organically depending on the distribution and number of instances observed at each leaf of the tree:
    • for nodes with low activity, DFDT deactivates the leaf node, conserving memory and computational resources.
    • for nodes with moderate activity, DFDT applies conservative splitting constraints based on entropy, information gain, and the number of instances observed at the leaf, ensuring controlled growth.
    • for highly active nodes, DFDT potentially employs a skipping mechanism, enabling rapid growth in response to significant data changes.
  • dynamically adjustable grace periods and tie thresholds, allowing the algorithm to either delay or accelerate splits depending on data variability. This flexibility improves the accuracy–memory–runtime trade-off and reduces the need for extensive hyperparameter tuning.
An ablation study of these DFDT’s components reveals three competitive variants tailored to different accuracy–memory trade-offs, summarized in Figure 1. These variants are fully compatible with existing ensemble frameworks (gomes2017survey; krawczyk2017ensemble), offering a drop-in alternative to standard VFDT-based learners.
Related Work
@@
The core component for incrementally constructing decision trees (DTs) on edge devices, while maintaining a fixed time complexity per sample, is approximation-based splitting (domingos2000mining). As new instances arrive, they are processed from the root to a leaf node, updating statistics at each node. These updates are used to periodically adjust heuristic values for each attribute, such as information gain (IG) and Gini index (GI) (domingos2000mining). The tree attempts to perform splits based on a statistical bound that quantifies the confidence interval for the heuristic function, given a minimum amount of data. Typically, DTs compare the Hoeffding bound (HB) to the difference in evaluation between the best and second-best attribute splits (domingos2000mining). When this difference exceeds the bound, the leaf node is split into child nodes.
Splitting rules.
@@
While traditional incremental DTs exclusively focus on evaluating the top two attributes, one can introduce extra splitting rules to promote even more valuable splits. For example, M-VFDT (yang2011moderated) incorporates the fluctuation of the HB, tracking its mean, minimum, and maximum values, along with an accuracy metric, as a pre-pruning condition for split decisions. Alternatively, SVFDT (da2018strict) applies constraints that compare current metrics against historical data and cross-leaf information, while introducing a skipping mechanism to bypass these constraints when significant changes in the data are detected. Furthermore, REG-VFDT (barddal2020regularized) adds a constraint to assess whether the best-ranked feature, chosen for splitting at a leaf node, provides substantial gains compared to the gains observed in previous splits within the same branch.
Tie breaking threshold ($\tau$).
@@
A critical aspect of incremental DTs is the tie-breaking procedure. While using the statistical difference in IG between two attributes helps control tree growth, competition between two equally favored split candidates can hinder progress, especially when either option would be equally suitable. To mitigate this, if the difference in heuristic values exceeds a predefined tie threshold, denoted as $\tau$, the split is performed. This threshold effectively controls the minimum rate of tree growth, with the attributes’ ability to separate before reaching the threshold influencing the speed at which the tree expands. However, a fixed $\tau$ value may cause ties to be broken prematurely, before a meaningful decision can be made, due to a lack of suitable candidates rather than a true tie situation. To address this, VFDT-$\tau$ (holmes2005tie) compares the difference between the best and second-best candidates with the difference between the second-best and worst candidates, while M-VFDT (yang2011moderated) designs an adaptive tie threshold that is dynamically calculated from the mean of HB, which has been shown to be proportionally related to the input stream samples.
Grace period ($n_{\min}$).
@@
To prevent premature splits with small sample sizes that undermine the validity of the HB, a grace period $n_{\min}$ specifies the minimum number of instances a leaf must observe before considering a split. However, without leveraging any information from the processed data, this can still result in computationally expensive split attempts or unnecessary delays in predictions. An alternative approach involves detecting frequent tie-breaking situations and applying steps to reduce the frequency of ties. For example, the default tie-breaking wait period can be adjusted by increasing the wait period for the child nodes after each tie is broken. This effect is cumulative until a valid split is found, after which the wait period is reset to the default value (holmes2005tie). Additionally, local statistics can be used to predict the optimal split time, minimizing delays and unnecessary split attempts. For instance, OSM (losing2018enhancing) uses class distributions from previous split attempts to estimate the minimum number of examples required before the HB. Alternatively, VFDT-$n_{\min}$ (garcia2018hoeffding) adjusts the grace period after unsuccessful split attempts, according to the reason of failure in order to ensure a split in the next iteration. For example, if the best attributes are not too similar, but their IG difference is insufficient to trigger a split, the solution is to wait for additional examples until the HB decreases enough to be smaller than the IG difference, adjusting the $n_{\min}$ accordingly. Similarly, if the top attributes are very similar in terms of IG, but the tie threshold $\tau$ is not exceeded, more instances are needed to allow the HB to decrease below $\tau$, with $n_{\min}$ adjusted based on their IG difference (garcia2018hoeffding).
Leaf activity.
@@
Another memory-efficient strategy is to condition split decisions based on leaf activity. GAHT (garcia2022green) devises a normalized measure to quantify how many instances have been observed at a particular leaf relative to the mean number of instances per leaf since its creation. Based on this value, nodes can either be deactivated when their activity is low, halting further splits to conserve resources, or more aggressive expansion strategies can be applied when activity exceeds a threshold, accelerating the growth of important nodes by relaxing the splitting rule to a less strict alternative. These adaptive expansion modes offer a more nuanced approach to tree growth, ensuring computational resources are focused on expanding nodes that significantly influence the model’s accuracy, while avoiding unnecessary splits in less relevant parts of the tree. Alternatively, VFDT (domingos2000mining) deactivates the least promising leaves based on the probability that examples will reach those leaves and their observed error rate. If a better split is found, the leaves of that split are deactivated, and their statistics are preserved. A new split is then performed, creating new leaves. If, during split re-evaluation, a previously deactivated split is found to be the best option, the saved statistics are restored rather than starting the process from scratch.
Overview. Table 1 summarizes how various algorithms implement different pre-pruning strategies, categorized by whether they adapt rules, grace periods, tie thresholds, or activity-based controls. As shown, DFDT offers the most comprehensive integration of these mechanisms, seamlessly combining them in a cohesive manner, as will be discussed.
Table 1: Incremental decision trees
Year Model Activity Rules $n_{\text{min}}$ $\tau$
2000 VFDT X
2005 VFDT-$\tau$ X X
2011 M-VFDT X X
2018 SVFDT X
2018 OSM X
2018 VFDT-$n_{\text{min}}$ X
2020 REG-VFDT X
2022 GAHT X
2026 DFDT X X X X
Methodology
@@
DFDT modifications are here described in detail, while referring to the pseudo-code in Algorithm 1. The initialization phase creates the root node structure of the decision tree (Alg. 1, Step 2, 3), sets up estimators (Alg. 1, Step 4), and instance counters (Alg. 1, Step 5). This initialization involves only constant-time operations, resulting in a computational complexity of $O(1)$, independent of the number of instances or features. The main loop iterates over each instance $(X,y)$ in the data stream $S$, running $N$ times in total. For each instance, several steps are performed. First, the instance is routed through the tree to the corresponding leaf node, where the prediction $\hat{y}$ is obtained (Alg. 1, Step 7). Assuming a balanced tree, the depth of traversal is $O(\log_{B}|LH|)$, where $|LH|$ is the number of leaf nodes and $B$ is the branching factor. The prediction at the leaf node is a constant-time operation, $O(1)$. Following this, the instance count and feature estimators are updated at the leaf node (Alg. 1, Steps 8-10). This update takes $O(F)$ time.
@@
1:procedure DFDT($S$: data stream, $\delta$: confidence level, $f_{\text{expand}}$: threshold for skipping, $f_{\text{deactivate}}$: threshold to deactivate leaves)
2:  $DFDT\leftarrow$ root
3:  $LH\leftarrow\text{hash of leaves}$
4:  Initialize $H_{LH_{\text{stat}}},n_{\text{stat}},H_{\text{stat}},G_{\text{stat}},HB_{\text{stat}},n_{\text{min}}$
5:  $n,n_{\text{l}},n_{\text{check}_{\text{l}}},n_{\text{leaf}_{\text{l}}},n_{\text{tree}_{\text{l}}}\leftarrow 0$ where $l=\text{root}$
6:  for each $(X,y)$ in $S$ do
7:   Route $(X,y)$ to leaf $l$ and obtain prediction $\hat{y}$
8:   Update $l$ feature estimators and class distribution
9:   $n\leftarrow n+1$
10:   $n_{l}\leftarrow n_{l}+1$
11:   if $\frac{(n_{l}-n_{\text{leaf}_{l}})\times|LH|}{n-n_{\text{tree}_{l}}}<f_{\text{deactivate}}$ then
12:     Deactivate $l$
13:   else if $f_{\text{expand}}$ then
14:     GrowFast $\leftarrow$ True
15:   end if
16:   if (impure at $l$) and ($n_{l}-n_{\text{check}_{l}}>n_{\text{min}}$) then
17:     Update $HB_{stat}$ with $\epsilon$
18:     if CanSplit(GrowFast) then
19:      Replace leaf $l$ with a split node
20:      for each leaf branch $b$ of the split do
21:        Initialize estimators
22:        Set post-split distribution at $b$
23:        Update LHLH
24:        $n_{b},n_{\text{leaf}_{b}},n_{\text{check}_{b}}\leftarrow\sum_{c\in\mathcal{C}}n_{b}^{(c)}$
25:        $n_{\text{tree}_{b}}\leftarrow n$
26:      end for
27:     else
28:      $n_{\text{check}_{l}}\leftarrow n_{l}$
29:      if $\overline{HB}_{\text{stat}}<\Delta G<\epsilon$ then
@@
30:        $n_{\text{min}}\leftarrow\left\lceil\frac{R^{2}\ln(1/\delta)}{2(\Delta G)^{2}}\right\rceil$
31:      else if $\Delta G<\overline{HB}_{\text{stat}}<\epsilon$ then
@@
32:        $n_{\text{min}}\leftarrow\left\lceil\frac{R^{2}\ln(1/\delta)}{2(\overline{HB}_{\text{stat}})^{2}}\right\rceil$
33:      end if
34:     end if
35:   end if
36:  end for
37:end procedure
@@
While the HB ensures that splits are made with statistical confidence, it treats all nodes equally in terms of expansion, despite the fact that not all nodes contribute equally to the decision-making process. To improve VFDT’s efficiency in allocating computational resources, a notion of relative importance can be introduced to dynamically adjust the rate of expansion at each node. To assess the activity level at each node, a fraction parameter is calculated as:
$\text{f}=\frac{(n_{l}-n_{\text{leaf}_{l}})\times|LH|}{n-n_{\text{tree}_{l}}}$ (1)
@@
where $(n_{l}-n_{\text{leaf}_{l}})$ is the number of instances observed at a particular leaf $l$ since its creation, $n-n_{\text{tree}_{l}}$ the total instances observed by the tree since the creation of leaf $l$, and $|LH|$ the current total number of leaves. If the fraction is below a threshold $f_{\text{deactivate}}$, the leaf node is deactivated, halting further splits to save on resources (Alg. 1, Steps 12), which is a constant-time operation, $O(1)$. If the metric falls above a threshold $f_{\text{expand}}$, a boolean saves the intention to apply a more aggressive expansion strategy, accelerating the growth of important nodes (Alg. 1, Step 14).
@@
To avoid costly evaluations, the algorithm only checks for potential splits once a leaf node accumulates $n_{\text{min}}$ instances since the last splitting opportunity (Alg. 1, Step 16). Then a split attempt is performed (Alg. 1, Step 18), with its pseudocode detailed in Algorithm 2. This check involves computing $G(\cdot)$ values, requiring $O(F)$ time, and sorting the $G(\cdot)$ values (Alg. 2, Step 2), yielding a time complexity of $O(F\log F)$. Next, the algorithm verifies whether the HB or tie-breaking threshold is satisfied (Alg. 2, Step 4). While in the VFDT, the tie threshold ($\tau$), which effectively controls minimum growth speed, remains static, DFDT, instead, dynamically sets the tie threshold as the mean of the Hoeffding Bound values computed over the most recent $k$ instances (yang2011moderated). This reduces the need for trial-and-error in selecting a fixed threshold, optimizing tree performance in dynamic environments. As the data stream becomes noisier, the adaptive $\tau$ stabilizes the model. If the HB is satisfied, DFDT utilizes adaptive expansion modes, determined by the fraction-determined boolean. For nodes with moderate activity, DFDT applies four conservative splitting constraints (da2018strict), ensuring controlled growth (Alg. 2, Step 12-15). For highly active nodes, i.e., when the fraction value exceeds 2, DFDT also checks two skipping conditions (Alg. 2, Step 5-11).
@@
1:procedure CanSplit(GrowFast = False),
2:  $G_{\text{best}}=\max(SortedG(\cdot))$
3:  $G_{\text{second best}}=\max(SortedG(\cdot)\setminus\{G_{\text{best}}\})$
4:  if $G_{\text{best}}-G_{\text{second best}}\geq\epsilon$ or $\epsilon<\overline{HB}_{\text{stat}}$ then
@@
6:     $C1\leftarrow H_{l}\geq\overline{H}_{\text{stat}}+\sigma(H_{\text{stat}})$
7:     $C2\leftarrow G_{\text{best}}\geq\overline{G}_{\text{stat}}+\sigma(G_{\text{stat}})$
8:     if C1 andand C2 then